# arXiv:2406.03010v2[quant-ph] 02 Mar 2026

### v2

@@ II.1 Approximation with MPS
The state of $N$ qubits in quantum computers can be described with the SV as
@@
where $|\psi\rangle$, $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$, and $\sigma_{i}\in\{0,1\}$ denote the state of $N$ qubits in quantum computers, SV, and the physical index for the state of $i$-th qubit [3], respectively. It is difficult to deal with the SV, $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$, exactly due to the numerous $(2^{N})$ elements in the tensor.
To reduce the number of independent elements, we employ MPS as shown in FIG. 1 to approximate the rank-$N$ tensor $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$ using a set of rank-3 local tensors $\{M^{\sigma_{i}}\}$. The approximation of SV by MPS is expressed as
@@
Each local tensor $M^{\sigma_{i}}$ has $2D^{2}$ elements, which is assumed not to depend on $N$, enabling the simulation of quantum computers with a reduced number of elements. Unless otherwise stated, we will employ MPS to describe quantum states throughout this paper.
@@ II.2 Tensor updates in MPS
STEP2: Decompose the matrix $\tilde{T}_{(a_{i-1},\sigma_{i}),(a_{i+1},\sigma_{i+1})}$ into $\sum_{k=1}^{D^{\prime}}A_{a_{i-1},k}^{\sigma_{i}}S_{k,k}B_{k,a_{i+1}}^{\sigma_{i+1}}$ using SVD, where $A$ and $B$ are tensors reshaped from unitary matrices and $S$ is a diagonal matrix with singular values.
STEP3: Contract tensors between $A$ and $S$ or $B$ and $S$ to make two rank-3 tensors $\sum_{a_{i}}^{D^{\prime}}M_{a_{i-1},a_{i}}^{\sigma_{i}}M_{a_{i},a_{i+1}}^{\sigma_{i+1}}$.
@@ II.3 Canonical form
where $\boldsymbol{\sigma}$ denotes $\{\sigma_{1}\sigma_{2}\ldots\sigma_{N}\}$, $A^{\sigma_{l}}=A_{a_{l-1},a_{l}}^{\sigma_{l}}$ and $B^{\sigma_{l}}=B_{a_{l-1},a_{l}}^{\sigma_{l}}$ satisfy canonical conditions, in other words, $\sum_{a_{l-1},\sigma_{l}}A_{a_{l-1},a^{\prime}_{l}}^{*\sigma_{l}}A_{a_{l-1},a_{l}}^{\sigma_{l}}$ and $\sum_{a_{l},\sigma_{l}}B_{a_{l-1},a_{l}}^{\sigma_{l}}B_{a^{\prime}_{l-1},a_{l}}^{*\sigma_{l}}$ are equal to the identity matrix, $m$ denotes the position of the Schmidt coefficients, and $\Lambda$ denotes a canonical center [25]. Eq. (8) is also called mixed canonical form. Matrix form is also utilized, e.g., $\sum_{\sigma_{l}}A^{\sigma_{l}{\dagger}}A^{\sigma_{l}}=I$, where $\dagger$} denotes Hermitian conjugate and $I$ is the identity matrix. In CF, the local matrices with a physical index in subsystem X are transformed into left normalized matrices $\{A^{\sigma_{l}}\}$ and those in subsystem Y are transformed into right normallized matrices $\{B^{\sigma_{l}}\}$. $\left(A^{\sigma_{1}}\ldots A^{\sigma_{m-1}}\right)$ in Eq. (8) is obtained from the iterative operations of the QR decomposition $M_{(\sigma_{i},a_{l-1}),a_{l}}=\sum_{a^{\prime}_{l}}Q_{(\sigma_{i},a_{l-1}),a^{\prime}_{l}}R^{i}_{a^{\prime}_{l},a_{l}}(i=1,\cdots,m-1)$ and the contraction $R^{i}M^{\sigma_{i+1}}=\tilde{M}^{\sigma_{i+1}}(i=1,\cdots,m-2)$, where $M_{(\sigma_{i},a_{l-1}),a_{l}}$ is the matrix reshaped from $M^{\sigma_{i}}$, $Q^{i}$ and $R^{i}$ are the matrices given by the QR decomposition from $M^{\sigma_{i}}$, $A^{\sigma_{i}}$ is the tensor reshaped from $Q^{i}$, and a tensor with denotes the updated tensor in a operation. is abbreviated in the next operation. $\left(\tilde{B}^{\sigma_{m+1}}B^{\sigma_{m+2}}\ldots B^{N}\right)$ is also obtained from the iterative operations of the QR decomposition $M_{a_{l-1},(\sigma_{i},a_{l})}=\sum_{a^{\prime}_{l-1}}R^{i}_{a_{l-1},a^{\prime}_{l-1}}Q^{i}_{a^{\prime}_{l-1},(\sigma_{i},a_{l})}(i=N,N-1,\cdots,m+1)$ and the contraction $M^{\sigma_{i-1}}R^{i}=\tilde{M}^{\sigma_{i-1}}(i=N,N-1,\cdots,m+2)$, where $M_{a_{l-1},(\sigma_{i},a_{l})}$ is the matrix reshaped from $M^{\sigma_{i}}$, $Q^{i}$ and $R^{i}$ are the matrices given by the QR decomposition from $M^{\sigma_{i}}$, and $B^{\sigma_{i}}$ is the tensor reshaped from $Q^{i}$. $A^{\sigma_{m}}$ and $\Lambda$ are obtained from SVD of $\tilde{M}^{\sigma_{m}}=R^{m-1}M^{\sigma_{m}}R^{m+1}$, i.e., $\tilde{M}_{(\sigma_{m},a_{m-1}),a_{m}}=\sum_{a^{\prime}_{m}}U_{(\sigma_{m},a_{m-1}),a^{\prime}_{m}}\Lambda_{a^{\prime}_{m},a^{\prime}_{m}}V^{*}_{a_{m},a^{\prime}_{m}}$, where $\tilde{M}_{(\sigma_{m},a_{m-1}),a_{m}}$ is the matrix reshaped from $\tilde{M}^{\sigma_{m}}$, $U$ and $V$ are the unitary matrices and $\Lambda$ is the diagonal matrix given by the SVD decomposition from $\tilde{M}^{\sigma_{m}}$, and $A^{\sigma_{m}}$ is the tensor reshaped from $U$. $V^{{\dagger}}$ is contracted with $B^{\sigma_{m+1}}$. CF of MPS corresponds to Schmidt decomposition [25], i.e., $\left(A^{\sigma_{1}}\ldots A^{\sigma_{m}}\right)$, $\left(B^{\sigma_{m+1}}\ldots B^{N}\right)$ and $\Lambda$ correspond to Schmidt bases ($x_{k}$, and $y_{k}$) and Schmidt coefficients $S_{k,k}$, respectively. Note that, this paper employs S as singular values of arbitrary matrix regardless whether these are Schmidt coefficients, while $\Lambda$ as Schmidt coefficients. In other words, singular values of SV or the canonical center are described with either $S$ or $\Lambda$, while singular values of other matrices are only denoted by $S$. In CF-based quantum computer simulations, the truncation of the diagonal matrix $\Lambda$ in CF is equivalent to the truncation of $S$ in Eq. (7) given by rank-$N$ tensor $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$.
@@ III.1 Confirming Reduction in Complexity
Two truncation parameters were employed: the number of singular values $\leq 5$, and the singular values are truncated when the ratio relative to largest singular value is less than $0.0001$. The number of two-qubit gates $N$ equals the number of qubits, where $N$ is a parameter.
@@
FIG. 5 shows the runtime for quantum circuits. In detail, the runtime refers to the duration required to update MPS using given tensor-form representation of the quantum circuit, excluding the time required to compute physical observables from the MPS. The SU significantly reduced the runtime in the cases of a large number of qubits $N$, e.g., SU is approximately $230$ times faster than CF for $N=2000$. In a log scale view, the slope of the runtime against $N$ is approximately 2 for CF and 1 for SU, respectively. Roughly speaking, the runtime is proportional to $N^{2}$ for CF and proportional to $N$ for SU, thus SU reduces the computational costs by $O(N)$. This reduction in computational costs is understood as follows. The average distance between two time-consective two-qubit gates is $O(N)$ for randomly ordered two-qubit gates, which requires the change of position $m$ in Eq.(8) at $O(N)$ complexity. CF requires the $O(N)$ complexity given by the QR decomposition and absorbing the matrix $R$, but SU does not.
@@ III.2 Investigating Performance
Two truncation parameters were employed: the number of singular values $\leq 10$, and the singular values are truncated when the ratio relative to largest singular value is less than $0.0001$. The number of qubits $N$ is $15$. We examine $29$ random quantum circuits generated by Qiskit for each depth.

### v3

# Simplification of Tensor Updates Toward Performance-Complexity Balanced Quantum Computer Simulation

@@ II.1 Approximation with MPS
The state of NN qubits in quantum computers can be described with the SV as
@@
where $|\psi\rangle$, $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$, and $\sigma_{i}\in\{0,1\}$ denote the state of NN qubits in quantum computers, SV, and the physical index for the state of $i$-th qubit [3], respectively. It is difficult to deal with the SV, $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$, exactly due to the numerous $(2^{N})$ elements in the tensor.
To reduce the number of independent elements, we employ MPS as shown in FIG. 1 to approximate the rank-NN tensor $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$ using a set of rank-3 local tensors $\{M^{\sigma_{i}}\}$. The approximation of SV by MPS is expressed as
@@
Each local tensor $M^{\sigma_{i}}$ has $2D^{2}$ elements, which is assumed not to depend on NN, enabling the simulation of quantum computers with a reduced number of elements. Unless otherwise stated, we will employ MPS to describe quantum states throughout this paper.
@@ II.2 Tensor updates in MPS
STEP2: Decompose the matrix $\tilde{T}_{(a_{i-1},\sigma_{i}),(a_{i+1},\sigma_{i+1})}$ into $\sum_{k=1}^{D^{\prime}}A_{a_{i-1},k}^{\sigma_{i}}S_{k,k}B_{k,a_{i+1}}^{\sigma_{i+1}}$ using SVD, where AA and $B$ are tensors reshaped from unitary matrices and $S$ is a diagonal matrix with singular values.
STEP3: Contract tensors between AA and $S$ or $B$ and $S$ to make two rank-3 tensors $\sum_{a_{i}}^{D^{\prime}}M_{a_{i-1},a_{i}}^{\sigma_{i}}M_{a_{i},a_{i+1}}^{\sigma_{i+1}}$.
@@ II.3 Canonical form
where $\boldsymbol{\sigma}$ denotes $\{\sigma_{1}\sigma_{2}\ldots\sigma_{N}\}$, $A^{\sigma_{l}}=A_{a_{l-1},a_{l}}^{\sigma_{l}}$ and $B^{\sigma_{l}}=B_{a_{l-1},a_{l}}^{\sigma_{l}}$ satisfy canonical conditions, in other words, $\sum_{a_{l-1},\sigma_{l}}A_{a_{l-1},a^{\prime}_{l}}^{*\sigma_{l}}A_{a_{l-1},a_{l}}^{\sigma_{l}}$ and $\sum_{a_{l},\sigma_{l}}B_{a_{l-1},a_{l}}^{\sigma_{l}}B_{a^{\prime}_{l-1},a_{l}}^{*\sigma_{l}}$ are equal to the identity matrix, $m$ denotes the position of the Schmidt coefficients, and $\Lambda$ denotes a canonical center [25]. Eq. (8) is also called mixed canonical form. Matrix form is also utilized, e.g., $\sum_{\sigma_{l}}A^{\sigma_{l}{\dagger}}A^{\sigma_{l}}=I$, where $\dagger$} denotes Hermitian conjugate and II is the identity matrix. In CF, the local matrices with a physical index in subsystem X are transformed into left normalized matrices $\{A^{\sigma_{l}}\}$ and those in subsystem Y are transformed into right normallized matrices $\{B^{\sigma_{l}}\}$. $\left(A^{\sigma_{1}}\ldots A^{\sigma_{m-1}}\right)$ in Eq. (8) is obtained from the iterative operations of the QR decomposition $M_{(\sigma_{i},a_{l-1}),a_{l}}=\sum_{a^{\prime}_{l}}Q_{(\sigma_{i},a_{l-1}),a^{\prime}_{l}}R^{i}_{a^{\prime}_{l},a_{l}}(i=1,\cdots,m-1)$ and the contraction $R^{i}M^{\sigma_{i+1}}=\tilde{M}^{\sigma_{i+1}}(i=1,\cdots,m-2)$, where $M_{(\sigma_{i},a_{l-1}),a_{l}}$ is the matrix reshaped from $M^{\sigma_{i}}$, $Q^{i}$ and $R^{i}$ are the matrices given by the QR decomposition from $M^{\sigma_{i}}$, $A^{\sigma_{i}}$ is the tensor reshaped from $Q^{i}$, and a tensor with denotes the updated tensor in a operation. is abbreviated in the next operation. $\left(\tilde{B}^{\sigma_{m+1}}B^{\sigma_{m+2}}\ldots B^{N}\right)$ is also obtained from the iterative operations of the QR decomposition $M_{a_{l-1},(\sigma_{i},a_{l})}=\sum_{a^{\prime}_{l-1}}R^{i}_{a_{l-1},a^{\prime}_{l-1}}Q^{i}_{a^{\prime}_{l-1},(\sigma_{i},a_{l})}(i=N,N-1,\cdots,m+1)$ and the contraction $M^{\sigma_{i-1}}R^{i}=\tilde{M}^{\sigma_{i-1}}(i=N,N-1,\cdots,m+2)$, where $M_{a_{l-1},(\sigma_{i},a_{l})}$ is the matrix reshaped from $M^{\sigma_{i}}$, $Q^{i}$ and $R^{i}$ are the matrices given by the QR decomposition from $M^{\sigma_{i}}$, and $B^{\sigma_{i}}$ is the tensor reshaped from $Q^{i}$. $A^{\sigma_{m}}$ and $\Lambda$ are obtained from SVD of $\tilde{M}^{\sigma_{m}}=R^{m-1}M^{\sigma_{m}}R^{m+1}$, i.e., $\tilde{M}_{(\sigma_{m},a_{m-1}),a_{m}}=\sum_{a^{\prime}_{m}}U_{(\sigma_{m},a_{m-1}),a^{\prime}_{m}}\Lambda_{a^{\prime}_{m},a^{\prime}_{m}}V^{*}_{a_{m},a^{\prime}_{m}}$, where $\tilde{M}_{(\sigma_{m},a_{m-1}),a_{m}}$ is the matrix reshaped from $\tilde{M}^{\sigma_{m}}$, $U$ and $V$ are the unitary matrices and $\Lambda$ is the diagonal matrix given by the SVD decomposition from $\tilde{M}^{\sigma_{m}}$, and $A^{\sigma_{m}}$ is the tensor reshaped from $U$. $V^{{\dagger}}$ is contracted with $B^{\sigma_{m+1}}$. CF of MPS corresponds to Schmidt decomposition [25], i.e., $\left(A^{\sigma_{1}}\ldots A^{\sigma_{m}}\right)$, $\left(B^{\sigma_{m+1}}\ldots B^{N}\right)$ and $\Lambda$ correspond to Schmidt bases ($x_{k}$, and $y_{k}$) and Schmidt coefficients $S_{k,k}$, respectively. Note that, this paper employs S as singular values of arbitrary matrix regardless whether these are Schmidt coefficients, while $\Lambda$ as Schmidt coefficients. In other words, singular values of SV or the canonical center are described with either $S$ or $\Lambda$, while singular values of other matrices are only denoted by $S$. In CF-based quantum computer simulations, the truncation of the diagonal matrix $\Lambda$ in CF is equivalent to the truncation of $S$ in Eq. (7) given by rank-NN tensor $C^{\sigma_{1}\sigma_{2}\ldots\sigma_{N}}$.
@@ III.1 Confirming Reduction in Complexity
Two truncation parameters were employed: the number of singular values $\leq 5$, and the singular values are truncated when the ratio relative to largest singular value is less than $0.0001$. The number of two-qubit gates NN equals the number of qubits, where NN is a parameter.
@@
FIG. 5 shows the runtime for quantum circuits. In detail, the runtime refers to the duration required to update MPS using given tensor-form representation of the quantum circuit, excluding the time required to compute physical observables from the MPS. The SU significantly reduced the runtime in the cases of a large number of qubits NN, e.g., SU is approximately $230$ times faster than CF for $N=2000$. In a log scale view, the slope of the runtime against NN is approximately 2 for CF and 1 for SU, respectively. Roughly speaking, the runtime is proportional to $N^{2}$ for CF and proportional to NN for SU, thus SU reduces the computational costs by $O(N)$. This reduction in computational costs is understood as follows. The average distance between two time-consective two-qubit gates is $O(N)$ for randomly ordered two-qubit gates, which requires the change of position $m$ in Eq.(8) at $O(N)$ complexity. CF requires the $O(N)$ complexity given by the QR decomposition and absorbing the matrix $R$, but SU does not.
@@ III.2 Investigating Performance
Two truncation parameters were employed: the number of singular values $\leq 10$, and the singular values are truncated when the ratio relative to largest singular value is less than $0.0001$. The number of qubits NN is $15$. We examine $29$ random quantum circuits generated by Qiskit for each depth.